OpenStudy (anonymous):

solve for ''Y'' 2y=96

TheSmartOne (thesmartone):

divide by 2 on both sides..

TheSmartOne (thesmartone):

And bam you get y all by itself.

TheSmartOne (thesmartone):

@Markiplier You are seeing something different. I see 2y=96

TheSmartOne (thesmartone):

y=96/2 y=?

TheSmartOne (thesmartone):

y=48

OpenStudy (anonymous):

Divide both sides by 2 So 2y/2 = 96/2 Final Answer is 92/2 which equals 48

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

Welcome :)